[ A 3000 gallon vat contains 1400 gallons of an 8% [acid solution]. Bob is going to add x gallons of an 80% acid solution to vat. Write the concentration of the mixture as a function of x. Help???
Answer by TimothyLamb(4379)

volume of acid in the final mixture (gallons):
a = 0.08(1400) + 0.80x
a = 112 + 0.80x
---
concentration of acid in the final mixture (%):
c = ( a / (1400 + x) ) * 100
---
Answer:
concentration of acid in the final mixture (%):
c = ( (112 + 0.80x) / (1400 + x) ) * 100
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Convert fractions, decimals, and percents:
https://sooeet.com/math/fraction-decimal-percent.php
---
Calculate and graph the linear regression of any data set:
https://sooeet.com/math/linear-regression.php